# Whisper D-SGD: Correlated Noise Across Agents for Differentially Private Decentralized Learning

### v1

@@ Abstract
Decentralized learning enables distributed agents to train a shared machine learning model through local computation and peer-to-peer communication. Although each agent retains its dataset locally, the communication of local models can still expose private information to adversaries. To mitigate these threats, local differential privacy (LDP) injects independent noise per agent, but it suffers a larger utility gap than central differential privacy (CDP). We introduce Whisper D-SGD, a novel covariance-based approach that generates correlated privacy noise across agents, unifying several state-of-the-art methods as special cases. By leveraging network topology and mixing weights, Whisper D-SGD optimizes the noise covariance to achieve network-wide noise cancellation. Experimental results show that Whisper D-SGD cancels more noise than existing pairwise-correlation schemes, substantially narrowing the CDP-LDP gap and improving model performance under the same privacy guarantees.
@@ I Introduction
Training machine learning models traditionally involves centralizing data on a single server, which raises scalability and privacy risks [1]. To address these issues, federated learning allows distributed agents (e.g., personal devices, organization, or computing nodes in a network) to retain their data locally while sharing only data-dependent computations (e.g., local gradients or models) with a central server [2, 3]. Decentralized learning further removes the need for a central server by allowing agents to update local models and mix them directly with neighbors according to predefined mixing weights [4, 5, 6, 7]. Although federated and decentralized learning algorithms avoid sharing raw data over the network, local models may still expose sensitive information to adversaries–i.e., external eavesdroppers–through attacks, e.g., membership inference or gradient inversion [8].
A widely adopted theoretical framework for mitigating such threats is differential privacy (DP), which involves injecting random noise into data-dependent computations, under strong, formal privacy guarantees [9]. While the privacy-utility trade-off of DP has been extensively studied in the central DP (CDP) setting–where a server injects noise on the central model [10, 11, 12]–adapting DP to decentralized settings proves substantially more challenging. Under the local DP (LDP) setting [13, 14], each agent individually adds noise to its local model before mixing, protecting its local dataset from eavesdroppers. However, LDP demands substantially more noise than CDP to attain the same privacy guarantees [15].
Recent works on differentially private decentralized learning seek to reconcile the utility of CDP with the local privacy of LDP by combining (i) a small amount of independent noise per agent, sufficient to ensure local privacy, and (ii) a large amount of correlated noise among agents, designed to cancel out during mixing [16, 17, 18]. In principle, this cancellation preserves local privacy while reducing the overall noise after mixing, such that the learned models approach CDP-level utility. Among these works, [16, 17] rely on zero-sum correlation, which is well-suited for addressing honest-but-curious internal neighbors, but ineffective against external eavesdroppers as no privacy noise persists in the final model. Meanwhile, [18] proposes pairwise-canceling correlated noise, which restricts noise correlation to pairs of neighboring agents, overlooking how local models are mixed across the network and limiting the potential for broader correlation design.
In this paper, we argue that extending beyond pairwise, neighbor-only correlations can significantly improve the privacy-utility trade-off in decentralized learning. The key contributions of this paper are as follows:
@@
  • We generalize existing analyses on the privacy-utility tradeoff, showing that the optimal covariance matrix must account for the network topology and mixing weights. Building on this insight, we propose the Whisper D-SGD algorithm, which optimizes the noise covariance to reconcile CDP-level utility with LDP-level privacy.
Experiments with various privacy budgets and network connectivity levels show that Whisper D-SGD amplifies noise cancellation over pairwise approaches, achieving superior privacy-utility trade-offs, particularly in weakly connected networks.
@@ II Privacy-Preserving Decentralized Learning
We consider $n$ agents, $\mathcal{V}=\{1,\dots,n\}$, each with a local dataset $\mathcal{D}_{i}$, aiming to learn the parameters $x\in\mathbb{R}^{d}$ of a shared machine learning model, by minimizing the global objective:
$F(x)\triangleq\frac{1}{n}\sum_{i=1}^{n}\left[F_{i}(x)\triangleq\frac{1}{|\mathcal{D}_{i}|}\sum_{\xi_{i}\in\mathcal{D}_{i}}\ell(x,\xi_{i})\right],$ (1)
where $\ell(x,\xi_{i})$ is the loss of parameter $x$ on sample $\xi_{i}\in\mathcal{D}_{i}$ and $F_{i}(x)$ is the local objective, known only to agent $i$. Agents communicate over a network modeled by an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where an edge $(i,j)\in\mathcal{E}$ indicates that agents $i$ and $j$ are neighbors, i.e., they can directly communicate. For simplicity, we present the case of scalar parameters ($d=1$), with straightforward extension to vector parameters.
II-A Decentralized Stochastic Gradient Descent (D-SGD)
Problem (1) is commonly solved via decentralized optimization algorithms like Decentralized Stochastic Gradient Descent (D-SGD) [6]. Each iteration $t\in\{1,\dots,T\}$ involves two steps:
a) Local step: Each agent $i$ samples $\xi_{i}^{(t)}\sim\mathcal{D}_{i}$, computes stochastic gradient $g_{i}^{(t)}=\ell^{\prime}(x_{i}^{(t)},\xi_{i}^{(t)})$, and updates parameter:
@@ II-A Decentralized Stochastic Gradient Descent (D-SGD)
b) Mixing step: Agents exchange parameters with neighbors and compute a weighted average:
${x}_{i}^{(t+1)}=\sum_{j=1}^{n}w_{ij}x_{j}^{(t+\frac{1}{2})},$ (2b)
@@
where $w_{ij}\triangleq[\mathbf{W}]_{ij}$ are mixing weights defined by the mixing matrix $\mathbf{W}\in\mathbb{R}^{n\times n}$, with $w_{ij}=0$ whenever $(i,j)\notin\mathcal{E}$. Convergence conditions for D-SGD require $\mathbf{W}$ to be doubly stochastic ($\mathbf{W}\mathbf{1}=\mathbf{1}$ and $\mathbf{1}^{\top}\mathbf{W}=\mathbf{1}^{\top}$) and the second largest eigenvalue $\lambda_{2}(\mathbf{W}^{\top}\mathbf{W})$ strictly below one [19, Section II.B].
II-B Differentially Private D-SGD
The DP framework protects the final models $\{x_{i}^{(T)}\}_{i=1}^{n}$ against adversaries (e.g., external eavesdroppers). At each iteration $t$, agent $i$ clips its local gradient $g_{i}^{(t)}$ to norm $C$, then adds privacy noise $v_{i}^{(t)}$:
$\tilde{g}_{i}^{(t)}=\hat{g}_{i}^{(t)}+v_{i}^{(t)},\quad\textstyle\hat{g}_{i}^{(t)}=\min\left\{1,\frac{C}{|g_{i}^{(t)}|}\right\}\cdot g_{i}^{(t)}.$ (3)
The resulting update is $\tilde{x}_{i}^{(t+1)}=\sum_{j=1}^{n}w_{ij}(\tilde{x}_{j}^{(t)}-\eta_{t}\tilde{g}_{j}^{(t)})$.
By relating $v_{i}^{(t)}$ to privacy parameters $(\varepsilon,\delta)$, the DP framework guarantees that the final models $\{\tilde{x}_{i}^{(T)}\}_{i=1}^{n}$, trained on any adjacent datasets $\mathcal{D},\mathcal{D}^{\prime}$ (differing in at most one agent’s dataset) satisfy, for any measurable set $\mathcal{S}$ in the parameter space:
$\mathbb{P}\bigl{[}\tilde{x}_{i}^{(T)}(\mathcal{D})\in\mathcal{S}\bigr{]}\leq e^{\varepsilon}\mathbb{P}\bigl{[}\tilde{x}_{i}^{(T)}(\mathcal{D}^{\prime})\in\mathcal{S}\bigr{]}+\delta.$ (4)
Common designs for $v_{i}^{(t)}$ include:
  1. 1.
    LDP [13, 14]. This is the standard approach, with privacy noise $v_{i}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{ldp}}^{2})$ independent across agents;
  2. 2.
    DECOR [18]. The privacy noise is designed as $v_{i}^{(t)}=u_{i}^{(t)}+\sum_{j\in\mathcal{N}_{i}}v_{ij}^{(t)}$, where $u_{i}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{pair}}^{2})$ is independent local noise, and each pair of neighbors $(i,j)\in\mathcal{E}$ shares a correlated noise term $v_{ij}^{(t)}=-v_{ji}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{cor}}^{2})$ that cancel during mixing. Distinct pairs of neighbors produce mutually uncorrelated components, so that $v_{ij}^{(t)}$ is independent of $v_{k\ell}^{(t)}$ whenever $\{i,j\}\cap\{k,\ell\}=\emptyset$.
III A Covariance-Based Framework for
@@
Generating Correlated Noise Across Agents
@@
We reinterpret the differentially private D-SGD framework from a noise-covariance perspective. Instead of drawing noise components $\{v_{i}^{(t)}\}_{i=1}^{n}$ independently or from pairwise correlations, we allow agents to sample from a multivariate Gaussian:
@@
where $\mathbf{R}\in\mathbb{R}^{n\times n}$ is an arbitrary covariance matrix. By defining $\tilde{\mathbf{x}}^{(t)}\triangleq[\tilde{x}_{1}^{(t)},\dots,\tilde{x}_{n}^{(t)}]^{\top}$ and $\tilde{\mathbf{g}}^{(t)}=[\tilde{g}_{1}^{(t)},\ldots,\tilde{g}_{n}^{(t)}]^{\top}$, the perturbed model-update at iteration $t$ becomes:
@@
This covariance-based approach offers several advantages: (i) it allows for more flexible correlation structures with noise cancellation beyond pairs of neighbors; (ii) it is directly applicable to directed graphs, unlike pairwise correlation; and (iii) it recovers state-of-the-art approaches as special cases:
@@
Remark 1.
The privacy noise generation mechanism $\mathbf{v}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{R})$ can be implemented in a distributed fashion. A common practice is to share $\mathbf{R}$ and a random seed $s$ among the agents. At iteration $t$, each agent independently samples $\mathbf{s}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$ using the random seed $s$ and then locally computes $\mathbf{v}^{(t)}=\mathbf{R}^{1/2}\mathbf{s}^{(t)}$.
The covariance matrix $\mathbf{R}$ controls both the level of privacy (in terms of $(\varepsilon,\delta)$-DP guarantees) and the model-update variance (hence convergence speed or utility). Intuitively, the noise must be “sufficiently large” to ensure LDP-level privacy, yet “sufficiently small” to achieve CDP-level utility.
@@ IV Privacy-Utility Analysis
In this section, we study how the correlated noise in our general framework affects both convergence (utility) performance and $(\varepsilon,\delta)$-DP (privacy) level. Our analysis shows the interplay between noise covariance $\mathbf{R}$ and mixing matrix $\mathbf{W}$, yielding new insights into optimal noise design strategies.
@@ IV-A Utility Analysis
To analyze utility, we introduce a virtual sequence $\hat{\mathbf{x}}^{(t+1)}=\mathbf{W}(\tilde{\mathbf{x}}^{(t)}-\eta_{t}\hat{\mathbf{g}}^{(t)})$, which begins at the noisy iterate $\tilde{\mathbf{x}}^{(t)}$ but applies a noise-free D-SGD step. This construction isolates the error introduced by the privacy noise and simplifies previous analyses (e.g., [18, Theorem 2]).
@@
[…]
@@
Theorem 1 shows that each iteration of our covariance-based framework incurs a baseline D-SGD error (independent of $\mathbf{R}$) plus a variance term proportional to $\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})$. [^2] Due to space limits, we defer the full characterization of global convergence guarantees after $T$ iterations to an extended version of this paper.
@@
Another remark from Theorem 1 is that the variance $\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})$ is linear in $\mathbf{R}$: smaller $[\mathbf{R}]_{ii}$ components speed-up convergence. At extreme cases, one may try to minimize this variance by enforcing $\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})=0$:
(i) If $\mathbf{W}$ is invertible, then $\mathbf{R}=\mathbf{0}$ is the only solution, implying no added noise—and hence no privacy.
(ii) If $\mathbf{W}$ is singular, there exists a nontrivial null space: $\mathbf{W}\,\mathbf{z}=\mathbf{0}$, $\mathbf{z}\neq\mathbf{0}$. One can then construct $\mathbf{R}=\mathbf{z}\,\mathbf{z}^{\top}\neq\mathbf{0}$ such that $\mathbf{W}\mathbf{R}\mathbf{W}^{\top}=\mathbf{0}$; however, $\mathbf{R}$ becomes singular. Our privacy analysis will show that invertibility of $\mathbf{R}$ is necessary for general $(\varepsilon,\delta)$-DP guarantees.

### v2

# Optimizing Privacy-Utility Trade-off in Decentralized Learning with Generalized Correlated Noise

@@ Abstract
Decentralized learning enables distributed agents to collaboratively train a shared machine learning model without a central server, through local computation and peer-to-peer communication. Although each agent retains its dataset locally, sharing local models can still expose private information about the local training datasets to adversaries. To mitigate privacy attacks, a common strategy is to inject random artificial noise at each agent before exchanging local models between neighbors. However, this often leads to utility degradation due to the negative effects of cumulated artificial noise on the learning algorithm. In this work, we introduce CorN-DSGD, a novel covariance-based framework for generating correlated privacy noise across agents, which unifies several state-of-the-art methods as special cases. By leveraging network topology and mixing weights, CorN-DSGD optimizes the noise covariance to achieve network-wide noise cancellation. Experimental results show that CorN-DSGD cancels more noise than existing pairwise correlation schemes, improving model performance under formal privacy guarantees.
@@ I Introduction
Training machine learning models traditionally involves centralizing datasets on a single server, which raises scalability and privacy risks [1]. To address these issues, federated learning allows distributed agents to retain their data locally while sharing only data-dependent computations (e.g., local gradients or models) with a central server [2, 3]. Decentralized learning further removes the need for a central server by allowing agents to update their local models and mix them directly with neighbors according to predefined mixing weights [4, 5, 6, 7]. Although federated and decentralized learning algorithms avoid sharing raw data over the network, local models may still expose sensitive information to adversaries through membership inference or gradient inversion attacks [8].
A widely adopted theoretical concept for mitigating such threats is differential privacy (DP), which provides formal privacy guarantees by injecting random noise into data-dependent computations [9]. While the privacy-utility trade-off has been extensively studied in the central DP (CDP) setting–where a server injects noise on the central model [10, 11, 12]–adapting DP to decentralized settings proves substantially more challenging. Under the local DP (LDP) setting [13, 14], each agent individually adds noise to its local model before mixing, protecting its local dataset from adversaries. However, since each agent adds noise independently without coordination, the utility of LDP degrades in terms of model performance [15].
Recent works on differentially private decentralized learning aim to improve LDP utility by combining (i) a small amount of independent noise per agent, sufficient to ensure local privacy, and (ii) a large amount of correlated noise across agents, designed to cancel during mixing [16, 17, 18]. In principle, this cancellation design preserves local privacy while reducing the overall noise after mixing, thereby improving the accuracy of the learned model. Among these works, [16, 17] rely on zero-sum correlation, which is well-suited for addressing honest-but-curious internal neighbors, but ineffective against external eavesdroppers. Meanwhile, [18] proposes pairwise-canceling correlated noise, which restricts noise correlation to pairs of neighboring agents, overlooking how local models are mixed across the network and limiting the potential for broader correlation design.
In this paper, we argue that extending beyond pairwise, neighbor-only correlations can significantly improve the privacy-utility trade-off in decentralized learning. The key contributions are summarized as follows:
@@
  • We generalize existing analyses on the privacy-utility trade-off, showing that the optimal covariance matrix should account for the network topology and mixing weights. Building on this insight, we propose the CorN-DSGD algorithm, which optimizes the noise covariance to achieve network-wide noise cancellation.
Experiments with various privacy budgets and network connectivity levels show that CorN-DSGD achieves superior privacy-utility trade-offs compared to pairwise approaches, especially in weakly connected networks.
@@ II Privacy-Preserving Decentralized Learning
We consider $n$ agents, $\mathcal{V}=\{1,\dots,n\}$, each with a local dataset $D_{i}$, aiming to learn the parameters $x\in\mathbb{R}^{d}$ of a shared machine learning model, by minimizing the global objective:
$F(x)\triangleq\frac{1}{n}\sum_{i=1}^{n}\left[F_{i}(x)\triangleq\frac{1}{|D_{i}|}\sum_{\xi_{i}\in D_{i}}\ell(x,\xi_{i})\right],$ (1)
where $\ell(x,\xi_{i})$ is the loss of parameter $x$ on sample $\xi_{i}\in D_{i}$ and $F_{i}(x)$ is the local objective, known only to agent $i$. Agents communicate over a network modeled by an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where an edge $(i,j)\in\mathcal{E}$ indicates that agents $i$ and $j$ are neighbors, i.e., they can directly communicate. For simplicity, we present the case of scalar parameters ($d=1$), with straightforward extension to vector parameters.
II-A Decentralized Stochastic Gradient Descent (DSGD)
Problem (1) is commonly solved via decentralized optimization algorithms like Decentralized Stochastic Gradient Descent (DSGD) [6]. Each iteration $t\in\{1,\dots,T\}$ involves two steps:
a) Local step: Each agent $i$ computes a stochastic gradient $g_{i}^{(t)}=\ell^{\prime}(x_{i}^{(t)},\xi_{i}^{(t)})$, where $\xi_{i}^{(t)}$ is a data point sampled from agent $i$’s dataset $D_{i}$, and then updates its parameter:
@@ II-A Decentralized Stochastic Gradient Descent (DSGD)
b) Mixing step: Agents exchange parameters with neighbors over the network and compute a weighted average:
$\textstyle{x}_{i}^{(t+1)}=\sum_{j=1}^{n}w_{ij}x_{j}^{(t+\frac{1}{2})},$ (2b)
@@
where $w_{ij}\triangleq[\mathbf{W}]_{ij}$ are mixing weights defined by the mixing matrix $\mathbf{W}\in\mathbb{R}^{n\times n}$, with $w_{ij}=0$ whenever $(i,j)\notin\mathcal{E}$. Convergence conditions for DSGD require $\mathbf{W}$ to be doubly stochastic ($\mathbf{W}\mathbf{1}=\mathbf{1}$ and $\mathbf{1}^{\top}\mathbf{W}=\mathbf{1}^{\top}$) and the second largest eigenvalue $\lambda_{2}(\mathbf{W}^{\top}\mathbf{W})$ strictly below one [19, Section II.B].
II-B Differentially-Private DSGD
In decentralized learning, we aim to protect agents’ local models from potential adversaries, which can be either external eavesdroppers or curious internal agents. For clarity of exposition, we address external eavesdroppers in the remainder of the paper and defer the discussion on honest-but-curious agents to Section V-D. Our privacy goal is to prevent the adversary from inferring whether any specific agent—and by extension, its entire local dataset—is participating in the training process. [^1] An alternative, weaker privacy objective is to protect each individual data sample rather than the entire dataset of an agent. Our framework extends to this sample-level DP setting in a straightforward manner.
Two sets of datasets $D=\{D_{1},D_{2},\dots,D_{n}\}$ and $D^{\prime}=\{D^{\prime}_{1},D^{\prime}_{2},\dots,D^{\prime}_{n}\}$ are agent-level neighbors if they differ by exactly one agent’s data ($D_{j}\neq D^{\prime}_{j}$ for some $j$, and $D_{i}=D^{\prime}_{i}$ for all $i\neq j$). A randomized mechanism $M(\cdot)$ is agent-level $(\varepsilon,\delta)$-DP if, for all neighboring datasets $D,D^{\prime}$ and all subsets $S$ of outputs, we have:
$\mathrm{Pr}\bigl{[}M(D)\in S\bigr{]}\leq e^{\varepsilon}\mathrm{Pr}\bigl{[}M(D^{\prime})\in S\bigr{]}+\delta.$ (3)
In the absence of a central server, privacy guarantees must be enforced locally by each agent. A classical LDP mechanism is presented in [13]. At each iteration $t$, each agent clips its own local gradient: $\hat{g}_{i}^{(t)}=\min\{1,\frac{C}{|g_{i}^{(t)}|}\}\cdot g_{i}^{(t)}$. It then adds noise locally and independently: $\tilde{g}_{i}^{(t)}=\hat{g}_{i}^{(t)}+v_{i}^{(t)}$. Finally, it applies the mixing step: $\tilde{x}_{i}^{(t+1)}=\sum_{j=1}^{n}w_{ij}(\tilde{x}_{j}^{(t)}-\eta_{t}\tilde{g}_{j}^{(t)})$. Specifically, the LDP Gaussian mechanism adds noise $v_{i}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{ldp}}^{2})$, such that, for sufficiently large noise variance $\sigma_{\text{ldp}}^{2}$, all local models $\{\tilde{x}_{i}^{(t+1)}\}_{i=1}^{n}$ are agent-level $(\varepsilon,\delta)$-LDP [14].
Despite its strong privacy guarantees, this LDP mechanism suffers from utility degradation, since the added independent privacy noise accumulates over the network, leading to higher variance in the mixed models, slower convergence, and reduced model performance [15]. Recognizing the utility degradation inherent to LDP, prior work [18] proposed pairwise-canceling correlated noise across agents. In DECOR [18], each agent adds noise $v_{i}^{(t)}=u_{i}^{(t)}+\sum_{j\in\mathcal{N}_{i}}v_{ij}^{(t)}$, where $u_{i}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{pair}}^{2})$ is independent local noise, and each pair of neighbors $(i,j)\in\mathcal{E}$ shares a correlated noise term $v_{ij}^{(t)}=-v_{ji}^{(t)}\sim\mathcal{N}(0,\sigma_{\text{cor}}^{2})$ that cancels in the mixing step. Distinct pairs of neighbors produce mutually uncorrelated components, so that $v_{ij}^{(t)}$ is independent of $v_{k\ell}^{(t)}$ whenever $\{i,j\}\cap\{k,\ell\}=\emptyset$. However, DECOR restricts noise correlations to neighboring agents and offers fewer degrees of freedom in noise design, particularly in sparse networks.
This paper introduces a novel covariance-based framework for generating correlated noise across agents, showing how the correlation structure should be designed by taking into account the effects of network topology and mixing weights.
@@ II-B Differentially-Private DSGD
III A Covariance-Based Framework for Generating Correlated Noise Across Agents
@@
We reinterpret the differentially private DSGD framework from a noise-covariance perspective. Instead of drawing noise components $\{v_{i}^{(t)}\}_{i=1}^{n}$ independently or from pairwise correlations, we allow agents to sample from a multivariate Gaussian:
@@
where $\mathbf{R}\in\mathbb{R}^{n\times n}$ is an arbitrary covariance matrix. By defining $\tilde{\mathbf{x}}^{(t)}\triangleq[\tilde{x}_{1}^{(t)},\dots,\tilde{x}_{n}^{(t)}]^{\top}$ and $\tilde{\mathbf{g}}^{(t)}=[\tilde{g}_{1}^{(t)},\ldots,\tilde{g}_{n}^{(t)}]^{\top}$, the perturbed model update at iteration $t$ becomes:
@@
In practice, each agent can generate $\mathbf{v}^{(t)}$ locally by sharing $\mathbf{R}$ and a random seed $s$, drawing $\mathbf{s}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$ independently using the seed $s$, and locally computing $\mathbf{v}^{(t)}=\mathbf{R}^{1/2}\mathbf{s}^{(t)}$.
This covariance-based approach offers several advantages: (i) it allows for more flexible correlation structures with noise cancellation beyond immediate neighbors; (ii) it is directly applicable to directed graphs, unlike pairwise correlation; and (iii) it recovers state-of-the-art approaches as special cases:
@@
Constraining $\mathbf{R}$ to the graph Laplacian $\sigma_{\text{cor}}^{2}\mathbf{L}$ limits noise correlation, particularly in sparse graphs. For instance, in a star topology, $\mathbf{R}=\sigma_{\text{pair}}^{2}\mathbf{I}_{n}+\sigma_{\text{cor}}^{2}\mathbf{L}$ forces $n(n-1)-2(n-1)$ covariance entries to zero, and overlooks all leaf-to-leaf noise cancellations, which a general $\mathbf{R}\succeq 0$ can instead exploit.
@@ IV Privacy-Utility Analysis
The covariance matrix $\mathbf{R}$ affects both utility (i.e., convergence rate) and privacy (i.e., $(\varepsilon,\delta)$-LDP guarantees).
@@ IV-A Utility Analysis
We introduce a virtual sequence $\hat{\mathbf{x}}^{(t+1)}=\mathbf{W}(\tilde{\mathbf{x}}^{(t)}-\eta_{t}\hat{\mathbf{g}}^{(t)})$, which begins at the noisy iterate $\tilde{\mathbf{x}}^{(t)}$ but applies a noise-free DSGD step. This construction isolates the error introduced by the privacy noise and simplifies the analysis in [18, Thm. 2].
@@
$=\;\underbrace{\mathbb{E}\Bigl{[}\bigl{\|}\hat{\mathbf{x}}^{(t+1)}-\tilde{\mathbf{x}}^{(t)}\bigr{\|}^{2}\;\Big{|}\;\mathcal{F}^{(t)}\Bigr{]}}_{\begin{subarray}{c}\text{Noise-free DSGD error}\\ \text{with clipped gradients}\end{subarray}}\;+\;\underbrace{\eta_{t}^{2}\,\mathrm{Tr}\!\bigl{(}\mathbf{W}\mathbf{R}\mathbf{W}^{\top}\bigr{)}}_{\text{Privacy-noise variance}}.$ (6)
@@
Theorem 1 shows that each iteration of our covariance-based framework incurs a baseline DSGD error (independent of $\mathbf{R}$) plus a variance term proportional to $\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})$.
@@
Additionally, the variance $\mathrm{Tr}(\mathbf{W}\mathbf{R}\mathbf{W}^{\top})$ is linear in $\mathbf{R}$: smaller $[\mathbf{R}]_{ii}$ components speed-up convergence. This implies that the optimal correlation structure should depend on the underlying network topology and mixing weights.